OpenStudy (kidthatbro8):

MEDAL AND FAN FOR ANSWER ! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and The profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, …where x is the number of smart phone cases sold and y is the number of tablet cases sold. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (igreen):

First let's change the equation from standard form to slope-intercept form. Standard Form: \(Ax + Bx = C\) Slope-Intercept Form: \(y = mx + b\) \(2x + 3y = 1200\) Subtract 2x to both sides: \(3y = -2x + 1200\) Divide 3 to all terms: \(\dfrac{3y}{3} = \dfrac{-2}{3}x + \dfrac{1200}{3}\) Can you simplify that? @Kidthatbro8

OpenStudy (igreen):

\(y = -\dfrac{2}{3}x + (1200 \div 3)\) Just divide 1200 / 3. What do you get? @Kidthatbro8

OpenStudy (kidthatbro8):

it's 4000.

OpenStudy (igreen):

4000?

OpenStudy (igreen):

I think you made a typo, check again.

OpenStudy (kidthatbro8):

400, sorry ^^

OpenStudy (igreen):

Right, so that gives us: \(y = -\dfrac{2}{3}x + 400\) Now it's in slope intercept form: \(y = mx + b\) Where m = slope, and b = y-intercept. So what's the slope and y-intercept of our equation? \(y = -\dfrac{2}{3}x + 400\)

OpenStudy (kidthatbro8):

the y-intercept is 400, i think.

OpenStudy (igreen):

Yep, and what's the slope?

OpenStudy (kidthatbro8):

2/3.

OpenStudy (igreen):

Almost. It's -2/3.

OpenStudy (igreen):

Now let's find the x-intercept, by plugging in 0 for y: \(y = -\dfrac{2}{3}x + 400\) \(0 = -\dfrac{2}{3}x + 400\) Subtract 400 to both sides: \(-400 = -\dfrac{2}{3}x\) Multiply -3/2 to both sides: What's -3/2 * -400? @Kidthatbro8

OpenStudy (kidthatbro8):

600?

OpenStudy (igreen):

Right so we have our y-intercept(0, 400) and our x-intercept (600, 0). Now to graph it we can just draw a line through these two points.

OpenStudy (igreen):

That answers your question..

OpenStudy (kidthatbro8):

thank you !